# How Semilocal Are Semilocal Density Functional Approximations? —Tackling Self-Interaction Error in One-Electron Systems

### v1

@@ Abstract
Figure 1: Binding Energy Curve of $H_{2}^{+}$ calculated with HF and different exchange-only density functionals.
@@
$E_{xc}[n]=\int d^{3}\textbf{r}n\epsilon_{xc}(n,\nabla n,\nabla^{2}n,\tau),$ (2)
In local density approximation (LDA) Von Barth and Hedin (1972); Perdew and Wang (1992), the electron density, $n(\textbf{r})$ is the only ingredient. Generalized gradient approximations (GGA) Perdew et al. (1996); Becke (1988); Perdew et al. (1992) add the electron density gradient, $\nabla n$. Meta-GGAs Perdew et al. (1999); Tao et al. (2003); Sun et al. (2012, 2015); Furness et al. (2020); Mejia-Rodriguez and Trickey (2017); Mejía-Rodríguez and Trickey (2020); Kaplan and Perdew (2022) further include the Laplacian of electron density, $\nabla^{2}n$, and/or the kinetic energy density, $\tau=\frac{1}{2}\sum_{i=1}^{N_{\mathrm{occ}}}|\nabla\phi_{i}|^{2}$, where $\phi_{i}$ are Kohn-Sham orbitals. With more ingredients, semilocal DFAs can satisfy more exact constraints, which are the known properties of the exact exchange-correlation energy Sun et al. (2015); Kaplan et al. (2023).
@@
To address this question, we start with the $\nabla^{2}n$-dependence meta-GGA exchange energy of a spin-unpolarized density,
@@
where $\epsilon_{x}^{unif}=-(3/4\pi)(3\pi^{2}n)$ is the exchange energy per particle of a uniform electron gas. $F_{x}(s,q)$ is the exchange enhancement factor depending on two dimensionless ingredients, the reduced density gradient $s=|\nabla n|/[2(3\pi^{2})^{1/3}n^{4/3}]$ and the reduced density laplacian $q=\nabla^{2}n/[4(3\pi^{2})^{2/3}n^{5/3}]$. The exchange energy for any spin-polarized density can be found from the exchange energy of a spin-unpolarized density using the exact spin-scaling relationship Oliver and Perdew (1979). We don’t consider the correlation energy functional in this study as $\tau$-dependent metaGGAs can be made one-electron self-correlation free Tao et al. (2003); Sun et al. (2015).
@@
The appropriate norms used here to determine $a$, $b$, and $q_{0}(s)$ are H and Gaussian electron density, which have analytic expressions (see section IV of Supplementary Material). Figure 2 (a) shows the system- and spherically-averaged exchange hole Hou et al. (2024a, b) against the inter-electron distance, $u$ for H and Gaussian electron density along with $H_{2}^{+}$ at various bond lengths. In comparison to H, Gaussian electron density as well as $H_{2}^{+}$ with the bond length, $R\leq 1.058$Å have deeper exchange holes, while $H_{2}^{+}$ with the bond length, $R\geq 1.270$Å have shallower ones. It suggests that Gaussian electron density as well as $H_{2}^{+}$ with the bond length, $R\leq 1.058$Å are as qualified for being appropriate norms as H for semilocal density functionals. Here, in addition to H, we choose Gaussian electron density over $H_{2}^{+}$ as an appropriate norm because Gaussian electron density is analytically known and not related to bonding. From the comparison between the PBE exchange holes and the exact ones given in section V of Supplementary Material, it is obvious that PBE’s exchange holes are more short-ranged than the exact ones, and become too deep for large R. It is also interesting to note that although the exact exchange holes of $H_{2}^{+}$ at equilibrium ($R=1.058$) and Gaussian electron density are deeper than those of H, SCAN predicts an overly negative exchange energy for the former (see Figure 1) and a less negative one for the latter (see Table 1). This suggests that the tailored GGA (eq. 4) may be limited by its variables and could benefit from incorporating $q$ for improved accuracy.
@@
$q_{0}(s)=s^{2}\left[1-\frac{2}{3\ln{\left((6\pi)^{1/3}(1+s^{2})\right)}}\right]$ (8)
by turning $s$ in the denominator into $1+s^{2}$. $q_{0}(s)$ eliminates the singularity of $q_{H}(s)$ and approaches $q_{H}(s)$ asymptotically from above, as shown in Figure 3 (a). This choice of $q_{0}(s)$ provides the flexibility for $g(s,q)\to 1$, thereby restoring $F_{x}^{SCAN}(s)$ for H and Gaussian electron density where $q_{H/G}(s)-q_{0}(s)<0$. For $H_{2}^{+}$, as the bond length $R$ increases, $q$ generally becomes more positive (see section II of the Supplementary Material), and thus $F_{x}^{RS}(s,q)$ tends to decrease as $g(s,q)$ is inversely proportional to q, which is desirable for a less negative exchange energy.
@@
However, after R = 4 Å, the improvement starts to diminish and gets slightly worse than SCAN as shown in Figure 1. This is because the (s, q) points appearing between $q_{0}(s)$ and $q_{H}(s)$ lines for example in Figure 4(b) for $R=4.76$ have too large q (q > 6), associated with relatively small s and very low electron densities (see section II of Supplementary Materials), and become energetically irrelevant. The energetically relevant electron density concentrates around the atomic centers with (s, q) below $q_{H}(s)$. This results in more negative exchange energy for RS than for SCAN because $F_{x}^{RS}$ decays slower than $F_{x}^{SCAN}$ with $s$ below $q_{H}(s)$. This is also dictated by the exchange hole of stretched bonds being too shallow, as shown in Figure 2.

### v2

@@ Abstract
Figure 1: $H_{2}^{+}$ binding energy curves of HF and exchange-only density functionals evaluated on HF orbitals.
@@
$E_{xc}[n]=\int d^{3}\textbf{r}n\epsilon_{xc}(n,\nabla n,\nabla^{2}n,\tau).$ (2)
In local density approximation (LDA) Von Barth and Hedin (1972); Perdew and Wang (1992), the electron density, $n(\textbf{r})$ is the only ingredient. Generalized gradient approximations (GGA) Perdew et al. (1996); Becke (1988); Perdew et al. (1992) add the electron density gradient, $\nabla n$. Meta-GGAs Perdew et al. (1999); Tao et al. (2003); Sun et al. (2012, 2015); Furness et al. (2020); Mejia-Rodriguez and Trickey (2017); Mejía-Rodríguez and Trickey (2020); Kaplan and Perdew (2022); Lebeda et al. (2024) further include the Laplacian of electron density, $\nabla^{2}n$, and/or the kinetic energy density, $\tau=\frac{1}{2}\sum_{i=1}^{N_{\mathrm{occ}}}|\nabla\phi_{i}|^{2}$, where $\phi_{i}$ are Kohn-Sham orbitals. With more ingredients, semilocal DFAs can satisfy more exact constraints, which are the known properties of the exact exchange-correlation energy Sun et al. (2015); Kaplan et al. (2023).
@@
To address this question, we start with the $\nabla^{2}n$-dependent meta-GGA exchange energy of a spin-unpolarized density,
@@
where $\epsilon_{x}^{unif}=-(3/4\pi)(3\pi^{2}n)$ is the exchange energy per particle of a uniform electron gas. $F_{x}(s,q)$ is the exchange enhancement factor depending on two dimensionless ingredients, the reduced density gradient $s=|\nabla n|/[2(3\pi^{2})^{1/3}n^{4/3}]$ and the reduced electron density Laplacian $q=\nabla^{2}n/[4(3\pi^{2})^{2/3}n^{5/3}]$. The exchange energy for any spin-polarized density can be found from the exchange energy of a spin-unpolarized density using the exact spin-scaling relationship Oliver and Perdew (1979). We don’t consider the correlation energy functional in this study as $\tau$-dependent metaGGAs can be made one-electron self-correlation free Tao et al. (2003); Sun et al. (2015).
@@
The appropriate norms used here to determine $a$, $b$, and $q_{0}(s)$ are H and Gaussian electron density, which have analytic expressions (see section IV of Supplementary Material). Figure 2 (a) shows the system- and spherically-averaged exchange hole Hou et al. (2024a, b) against the inter-electron distance, $u$, for H and Gaussian electron density along with $H_{2}^{+}$ at various bond lengths. In comparison to H, Gaussian electron density as well as $H_{2}^{+}$ with the bond length, $R\leq 1.058$Å have deeper exchange holes, while $H_{2}^{+}$ with the bond length, $R\geq 1.270$Å have shallower ones. It suggests that Gaussian electron density is as qualified for being an appropriate norm as H for semilocal density functionals since Gaussian electron density is also analytically known and not related to bonding. From the comparison between the PBE exchange holes and the exact ones given in section V of Supplementary Material, it is obvious that PBE’s exchange holes are more short-ranged than the exact ones, and become too deep for large bond lengths R of $H_{2}^{+}$. It is also interesting to note that although the exact exchange holes of $H_{2}^{+}$ at equilibrium ($R=1.058$) and Gaussian electron density are deeper than those of H, SCAN predicts an overly negative exchange energy for the former (see Figure 1) and a less negative one for the latter (see Table 1). This suggests that the tailored GGA (eq. 4) may be limited by its variables and could benefit from incorporating $q$ for improved accuracy.
@@
$q_{0}(s)=s^{2}\left[1-\frac{2}{3\ln{\left((6\pi)^{1/3}(1+s^{2})^{1/2}\right)}}\right]$ (8)
by turning $s$ in the denominator into $\sqrt{1+s^{2}}$. $q_{0}(s)$ eliminates the singularity of $q_{H}(s)$ and approaches $q_{H}(s)$ asymptotically from above, as shown in Figure 3 (a). This choice of $q_{0}(s)$ provides the flexibility for $g(s,q)\to 1$, thereby restoring $F_{x}^{SCAN}(s)$ for H and Gaussian electron density where $q_{H/G}(s)-q_{0}(s)<0$. For $H_{2}^{+}$, as the bond length $R$ increases, $q$ generally becomes more positive (see section II of the Supplementary Material), and thus $F_{x}^{RS}(s,q)$ tends to decrease as $g(s,q)$ is inversely proportional to q, which is desirable for a less negative exchange energy.
@@
However, after R = 4 Å, the improvement starts to diminish and gets slightly worse than SCAN as shown in Figure 1. This is because the (s, q) points for $R=4.76$ appearing above $q_{0}(s)$ have too large q (q > 6) as shown in Figure 4(b), associated with relatively small s and very low electron densities (see section II of Supplementary Materials), and become energetically irrelevant. The energetically relevant electron density concentrates around the atomic centers with (s, q) below $q_{H}(s)$. This results in more negative exchange energy for RS than for SCAN because $F_{x}^{RS}$ decays slower than $F_{x}^{SCAN}$ with $s$ below $q_{H}(s)$. This is also dictated by the exchange hole of stretched bonds being too shallow, as shown in Figure 2.